# GraphCogent: Overcoming LLMs’ Working Memory Constraints via Multi-Agent Collaboration in Complex Graph Understanding

### v1

@@ Abstract
Large language models (LLMs) show promising performance on small-scale graph reasoning tasks but fail when handling real-world graphs with complex queries. This phenomenon stems from LLMs’ inability to effectively process complex graph topology and perform multi-step reasoning simultaneously. To address these limitations, we propose GraphCogent, a collaborative agent framework inspired by human Working Memory Model that decomposes graph reasoning into specialized cognitive processes: sense, buffer, and execute. The framework consists of three modules: Sensory Module standardizes diverse graph text representations via subgraph sampling, Buffer Module integrates and indexes graph data across multiple formats, and Execution Module combines tool calling and model generation for efficient reasoning. We also introduce Graph4real, a comprehensive benchmark contains with four domains of real-world graphs (Web, Social, Transportation, and Citation) to evaluate LLMs’ graph reasoning capabilities. Our Graph4real covers 21 different graph reasoning tasks, categorized into three types (Structural Querying, Algorithmic Reasoning, and Predictive Modeling tasks), with graph scales that are 10 times larger than existing benchmarks. Experiments show that Llama3.1-8B based GraphCogent achieves a 50% improvement over massive-scale LLMs like DeepSeek-R1 (671B). Compared to state-of-the-art agent-based baseline, our framework outperforms by 20% in accuracy while reducing token usage by 80% for in-toolset tasks and 30% for out-toolset tasks. Code will be available after review.
1 Introduction
@@
Large language models (LLMs) OpenAI (2023) have demonstrated remarkable cognitive capabilities in natural language processing. However, when querying the shortest path between two places in a 1000-node transportation graph, state-of-the-art LLMs like DeepSeek-R1 DeepSeek-AI (2025), GPT-o3 OpenAI (2023), and Gemini-2.5 pro Group (2023) make wrong turns in 9 out of 10 cases. This failure reveals that current LLMs remain limited in handling large real-world graph reasoning tasks.
Researchers have explored various methods to address the constraints of LLMs’ inability to perform graph reasoning. Text-based methods Fatemi et al. (2024); Luo et al. (2024); Chen et al. (2024) leverage Chain-of-Thought (CoT) Wei et al. (2022) reasoning but struggle with the dynamic programming requirements of graph algorithms (e.g., Bellman-Ford Bellman (1958)), resulting in cascading errors in complex tasks. Tool-based methods Zhang (2023); Wang et al. (2025b) rely on external tools for computation, requiring predefined tools and rigid input formats (e.g., preprocessed files). These limitations hinder their ability to process diverse graph text representations, such as adjacency lists, symbolic notations, and linguistic descriptions Fatemi et al. (2024), thereby reducing their adaptability to real-world scenarios. Agent-based methods Li et al. (2024) employ multi-agent collaboration, often decomposing graph tasks into sequential stages via naive instruction prompts. However, these approaches are fundamentally limited by the complexity of code construction and the inherent memory constraints of LLMs. As graph scales increase and tasks involve real-world graph problems, agent-based methods perform poorly when simultaneously handling graph data comprehension and task execution.
This limitation stems from LLMs’ working memory constraints that restrict simultaneous processing of complex graph topology and multi-step reasoning. A common mitigation approach involves partitioning graphs into smaller units for sequential processing Jiang et al. (2023), mirroring how humans decompose complex problems into manageable information chunks Miller (1956). Classic human working memory capacity assessments N-back test Kirchner (2006); Klatzky et al. (2008); Amon and Bertenthal (2018) (exampled in Fig. 1, left) reveals a rapid cognitive decline when humans are required to retain more than three items. Inspired by the N-back test, we design an analogous LLMs’ Graph N-back test. As exampled in the right of the Fig. 1, by the 3rd processing turn, Llama3.1-8B Touvron et al. (2023) exhibit significant topology forgetting, detailed quantitative result is analyzed in Section 3.1. This demonstrates that LLM graph reasoning similarly suffers from working memory constraints, mirroring human cognitive limitations.
@@
Figure 1: Graph N-back Query Task: A graph is split into 50-edge subsets $E_{t}$. At turn $t+N$, the LLM verifies edge existence in $E_{t}$. Experimental results are in Section 3.1.
@@
Human brain addresses these constraints through a specialized working memory model for sensory processing, information buffering, and task execution Baddeley et al. (2011). Facing a complex task, humans rely on an external sensory system to perceive stimuli, an episodic buffer to integrate and store information, and a central executive to coordinate cognition and processing. This absent architecture in monolithic LLMs reveals their key graph reasoning bottlenecks: (1) Diverse Text Representations: The sensory limitation in processing diverse graph text representations leads to inconsistent comprehension, especially for mainstream open-source models (e.g., Llama Touvron et al. (2023), Qwen Yang et al. (2024), and GLM GLM (2024)), thus impairing reasoning accuracy; (2) Overload Graph Scale: Lacking a buffer mechanism, LLMs’ limited input context windows hinders long-range dependency capture, causing global information loss and cascading reasoning errors; (3) Code Execution Fragility: The executive dysfunction results in unreliable code generation and inefficient algorithmic implementation.
@@
Inspired by the cognitive architecture of human working memory system, we propose GraphCogent, an innovative Graph Collaboration Agentic framework designed to overcome LLMs’ working memory constraints in graph reasoning. Our framework consists of three modules: Sensory Module employs a sensory agent to sample subgraph and transform unstructured graph text representations into standardized adjacency list. Buffer Module integrates graph data from the Sensory Module and establishes data indices for diverse data formats (including Numpy, PyG, and NetworkX) based on Buffer Agent. Execution Module synergistically combines tool calling and model generation. Naive graph reasoning tasks are handled by a Reasoning Agent through a pre-built common toolset, while complex tasks leverage a Model Generation Agent that produces modular, task-specific components, avoiding error-prone full code generation by building upon the Buffer’s preprocessed data.
@@
To evaluate LLM’s generalization capability in graph reasoning, we construct a more challenging Graph4real benchmark dataset, which collects graph data from four real-world domains: Web McAuley and Leskovec (2012), Social McAuley and Leskovec (2012), Transportation Shuman et al. (2013), and Citation Craven et al. (1998). Our Graph4real covers 21 different graph reasoning tasks, categorized into three types (Structural Querying, Algorithmic Reasoning, and Predictive Modeling tasks), with graph scales that are 10 times larger than existing benchmarks (e.g., NLGraph Wang et al. (2023), GraphWiz Chen et al. (2024)). Experimental results demonstrate that our method achieves state-of-the-art performance with an average accuracy of 98.5% based on the open-source Llama3.1-8B, representing a 20% improvement over other agent-based approaches. Cross-dataset validation demonstrates robust generalization, maintaining over 90% accuracy across diverse public benchmarks. The combining tool calling and model-generation strategy delivers dual optimization: it reduces token consumption by 80% for tasks covered by our toolset while still achieving 30% token savings for out-toolset tasks compared to the conventional agent-based method.
@@
Our contributions are summarized as follows: (1) We propose GraphCogent, a novel agent collaboration framework addressing LLMs’ working memory constraints through its sensory-buffer-execution architecture; (2) We develop a combining reasoning strategy that integrates tool calling and model generation to handle diverse graph tasks while optimizing computational efficiency; (3) We construct Graph4real, a comprehensive benchmark featuring real-world graphs that covers 21 different graph reasoning tasks, categorized into three types (Structural Querying, Algorithmic Reasoning, and Predictive Modeling tasks), with graph scales that are 10 times larger than existing benchmarks.
2 Graph4real Construction
@@
Real-World Graph Scaling. To establish a comprehensive benchmark for evaluating LLMs on real-world graph reasoning tasks, we curate a diverse set of graphs from four domains: Web (Google Web Graph McAuley and Leskovec (2012)), Social (SNAP Social Circles McAuley and Leskovec (2012)), Transportation (PeMS Shuman et al. (2013)), and Citation graphs (Cora Craven et al. (1998)). Unlike prior benchmarks Wang et al. (2023); Luo et al. (2024) that rely on randomly generated graphs, our dataset is constructed from real-world sources to ensure realistic topological properties and domain-specific characteristics. We employ a biased random walk sampling strategy to construct graphs at three scales (40, 100, and 1000 nodes), ensuring both effective evaluation of LLM-based methods through smaller-scale graphs and construct large-scale graphs to challenge existing methods.
@@
Intent-Driven Task Design. A total number of 21 tasks are designed and categorized into three classes: (1) Graph Structural Querying Tasks: Focus on fundamental graph properties, such as edge existence or node count, testing basic graph structural comprehension. (2) Graph Algorithmic Reasoning Tasks: Require reasoning based on classical algorithms, such as shortest path and maximum flow, evaluating algorithmic proficiency. (3) Graph Predictive Modeling Tasks: Involve neural network-based predictions, such as node classification and link prediction, assessing predictive modeling capabilities.
For task generation, we collect five real-world scenarios for each of the four domains (e.g., travel planning and logistics optimization for transportation) and craft 20 prompt templates to construct contextually grounded questions. Using DeepSeek-R1 DeepSeek-AI (2025), we generate graph reasoning tasks based on templates with specific scenarios, ensuring both diversity and relevance. To validate task quality, we employ a dual evaluation approach: DeepSeek-R1 for automated assessment and human for manual review. This process produces a dataset with 4200 questions distributed across the 21 tasks. The detailed task definitions, dataset statistics, and prompt templates are provided in Appendix F.
Figure 2: Overview of GraphCogent. Sensory Module (left) standardizes various graph text representations through subgraph sampling and conversion; Buffer Module (center) establishes cross-format data (e.g., NetworkX) integrating and indexing transformations; Execution Module (right) enables two reasoning modes: Reasoning Agent is employed for task discrimination and implements tool calling for in-toolset tasks, Model Agent handles out-toolset tasks based on model generation.
3 GraphCogent
We propose GraphCogent, an agent-based framework inspired by the human working memory system to address the challenges of real-world graph reasoning tasks. Our framework consists of three specialized modules: Sensory Module (samples subgraphs and transform unstructured graph text representations into standardized adjacency list), Buffer Module (integrates graph data from Sensory Module and establishes data indices for diverse data formats), and Execution Module (combines tool calling and model generation to comprehensively address graph reasoning tasks). The overall pipeline of the GraphCogent is illustrated in Fig. 2.
3.1 Sensory Module
Real-world graphs present two fundamental characteristics: (1) Large Graph Scale, benchmarks like NLGraph Wang et al. (2023) use small graphs (less than 40 nodes), while real-world graph (e.g., Cora Craven et al. (1998)) contains thousands of nodes and edges; (2) Varied Text Representation, different domains use distinct descriptions ("connected" in transportation or "followed" in social), and same text representation may appear in varied formats (adjacency lists, symbolic notations, linguistic descriptions).
@@ 3.1 Sensory Module
Figure 3: Graph N-back Query. Accuracy (right y-axis) measures memory retention across dialogues, computed as sum of True edges (correctly identified existing edges) and False edges (correctly rejected non-existent edges) (left y-axis).
@@
Figure 4: Graph Efficiency Coefficient (right y-axis) across sampling granularity. Edit shows graph Edit distance (left y-axis), and Cost (left y-axis, scaled ×100) counts total LLM interactions. Lower GEC indicates better performance.
@@
We propose Graph N-back Query, a graph memory evaluation for LLM inspired by the human working memory capacity assessment through N-back tasks Kirchner (2006) (Detailed design of the Graph N-back Query and the working memory inspiration behind our framework are provided in Appendix A): Using a 100-node (sampled from PeMS datasets Shuman et al. (2013)) partitioned into 50-edge subsets $E_{t}$, we test memory retention across three text representations by querying edge existence at turn $t+N$. Experiments with Llama3.1-8B in Fig. 4 reveal significant information forgetting across dialogues, demonstrating that LLMs’ inherent memory mechanisms fail to reliably handle large-scale graph reasoning. This highlights that LLMs’ inherent context window limitations constrains their ability to maintain global graph information, rendering conventional step-by-step graph input approaches ineffective due to these working memory constraints.

### v2

# GraphCogent: Mitigating LLMs’ Working Memory Constraints via Multi-Agent Collaboration in Complex Graph Understanding

@@ Abstract.
Large language models (LLMs) show promising performance on small-scale graph reasoning tasks but fail when handling real-world graphs with complex queries. This phenomenon arises from LLMs’ working memory constraints, which result in their inability to retain long-range graph topology over extended contexts while sustaining coherent multi-step reasoning. However, real-world graphs are often structurally complex, such as Web, Transportation, Social, and Citation networks. To address these limitations, we propose GraphCogent, a collaborative agent framework inspired by human Working Memory Model that decomposes graph reasoning into specialized cognitive processes: sense, buffer, and execute. The framework consists of three modules: Sensory Module standardizes diverse graph text representations via subgraph sampling, Buffer Module integrates and indexes graph data across multiple formats, and Execution Module combines tool calling and tool creation for efficient reasoning. We also introduce Graph4real, a comprehensive benchmark that contains four domains of real-world graphs (Web, Transportation, Social, and Citation) to evaluate LLMs’ graph reasoning capabilities. Our Graph4real covers 21 different graph reasoning tasks, categorized into three types (Structural Querying, Algorithmic Reasoning, and Predictive Modeling tasks), with graph scales up to 10 times larger than existing benchmarks. Experiments show that Llama3.1-8B based GraphCogent achieves a 50% improvement over massive-scale LLMs like DeepSeek-R1 (671B). Compared to state-of-the-art agent-based baseline, our framework outperforms by 20% in accuracy while reducing token usage by 80% for in-toolset tasks and 30% for out-toolset tasks. Code will be available after review.
Large Language Model, Graph Algorithms, Multi-agent System
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; 2026; $X$, XXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Networks Network algorithmsccs: Computing methodologies Knowledge representation and reasoningccs: Mathematics of computing Graph algorithms
1. Introduction
@@
Large language models (LLMs) (OpenAI, 2023) have demonstrated remarkable cognitive capabilities in natural language processing. However, when querying the connection path between two webpages in a Web graph, state-of-the-art LLMs like DeepSeek-R1 (DeepSeek-AI, 2025), GPT-o3 (OpenAI, 2023), and Gemini-2.5 pro (Group, 2023) return incorrect navigation routes in 9 out of 10 cases. This failure reveals that current LLMs remain limited in handling large real-world graph reasoning tasks.
@@
Figure 1. Graph N-back Query Task: A graph is split into 50-edge subsets $E_{t}$. At turn $t+N$, the LLM verifies edge existence in $E_{t}$. Experimental results are in Section 4.2.
Researchers have explored various methods to address the constraints of LLMs’ inability to perform graph reasoning. Text-based methods (Fatemi et al., 2024; Luo et al., 2024; Chen et al., 2024) leverage Chain-of-Thought (CoT) (Wei et al., 2022) reasoning but struggle with the dynamic programming requirements of graph algorithms (e.g., Bellman-Ford (Bellman, 1958)), resulting in cascading errors in complex tasks. Tool-based methods (Zhang, 2023; Wang et al., 2025) rely on external tools for computation, requiring predefined tools and rigid input formats (e.g., preprocessed files). These limitations hinder their ability to process diverse graph text representations, such as adjacency lists, symbolic notations, and linguistic descriptions (Fatemi et al., 2024), thereby reducing their adaptability to real-world scenarios. Agent-based methods (Li et al., 2024; Zhang et al., 2024) employ multi-agent collaboration, often decomposing graph tasks into sequential stages via naive instruction prompts. However, these approaches are fundamentally limited by the complexity of code construction and the inherent memory constraints of LLMs. As graph scales increase and tasks involve real-world graph problems, agent-based methods perform poorly when handling graph data comprehension and task execution.
This limitation stems from LLMs’ working memory constraints that restrict simultaneous processing of complex graph topology and multi-step reasoning. A common mitigation approach involves partitioning graphs into smaller units for sequential processing (Jiang et al., 2023), mirroring how humans decompose complex problems into manageable information chunks (Miller, 1956). Classic human working memory capacity assessments N-back test (Kirchner, 2006; Klatzky et al., 2008; Amon and Bertenthal, 2018) (exampled in Fig. 1, left) reveals a rapid cognitive decline when humans are required to retain more than three items. Inspired by the N-back test, we design an analogous LLMs’ Graph N-back test. As exampled in the right of the Fig. 1, by the 3rd processing turn, Llama3.1-8B (Touvron et al., 2023) exhibits significant topology forgetting, detailed quantitative result is analyzed in Section 4.2. This demonstrates that LLM graph reasoning similarly suffers from working memory constraints, mirroring human cognitive limitations.
@@
Human brain addresses these constraints through a specialized working memory model for sensory processing, information buffering, and task execution (Baddeley et al., 2011). Facing a complex task, humans rely on an external sensory system to perceive stimuli, an episodic buffer to integrate and store information, and a central executive to coordinate cognition and processing. This absent architecture in monolithic LLMs reveals their key graph reasoning bottlenecks: (1) Diverse Text Representations: The sensory limitation in processing diverse graph text representations leads to inconsistent comprehension, especially for mainstream open-source models (e.g., Llama (Touvron et al., 2023), Qwen (Yang et al., 2024), and GLM (GLM, 2024)), thus impairing reasoning accuracy; (2) Overload Graph Scale: Lacking a buffer mechanism, LLMs’ limited input context windows hinders long-range dependency capture, causing global information loss and cascading reasoning errors; (3) Code Execution Fragility: The executive dysfunction results in unreliable code generation and inefficient algorithmic implementation.
@@
Inspired by the cognitive architecture of human working memory model, we propose GraphCogent, a Graph Collaboration Agentic framework designed to mitigate LLMs’ working memory constraints in graph reasoning. Our framework consists of three modules: Sensory Module employs a sensory agent to sample subgraph and transform unstructured graph text representations into standardized adjacency list. Buffer Module integrates graph data from the Sensory Module and establishes data indices for diverse data formats (including NumPy, PyG, and NetworkX) based on Buffer Agent. Execution Module synergistically combines tool calling and with on-demand tool creation. Naive graph reasoning tasks are handled by a Execution Agent through a pre-built common toolset, while complex tasks leverage a tool creator that produces modular, task-specific components and backfills new tool requests with the Buffer’s preprocessed data, thereby avoiding error-prone full-code generation.
@@
To evaluate LLM’s generalization capability in graph reasoning, we construct a more challenging Graph4real benchmark dataset, which collects graph data from four real-world domains: Web (McAuley and Leskovec, 2012), Social (McAuley and Leskovec, 2012), Transportation (Shuman et al., 2013), and Citation (Craven et al., 1998). Our Graph4real covers 21 different graph reasoning tasks, categorized into three types (Structural Querying, Algorithmic Reasoning, and Predictive Modeling tasks), with graph scales that are 10 times larger than existing benchmarks (e.g., NLGraph (Wang et al., 2023), GraphWiz (Chen et al., 2024)). Experimental results demonstrate that our method achieves state-of-the-art performance with an average accuracy of 98.5% based on the open-source Llama3.1-8B, representing a 20% improvement over other agent-based approaches. Cross-dataset validation demonstrates robust generalization, maintaining over 90% accuracy across diverse public benchmarks. The combining tool calling and tool creation strategy delivers dual optimization: it reduces token consumption by 80% for tasks covered by our toolset while still achieving 30% token savings for out-toolset tasks compared to the conventional agent-based method.
@@
Our contributions are summarized as follows: (1) We propose GraphCogent, a novel agent collaboration framework addressing LLMs’ working memory constraints through its sensory-buffer-execution architecture; (2) We design a reasoning mechanism that combines tool calling and tool creation with a Thinking Policy Initialization to stabilize early-stage decision-making and a tailored Capability-Margin Preference Optimization (CMPO) objective to sharpen tool discrimination, enabling efficient handling of diverse graph tasks; (3) We construct Graph4real, a comprehensive benchmark featuring real-world graphs that covers 21 different graph reasoning tasks, categorized into three types (Structural Querying, Algorithmic Reasoning, and Predictive Modeling tasks), with graph scales that are 10 times larger than existing benchmarks.
2. Related work
LLMs for Graph Reasoning Tasks. Existing graph reasoning methods adopt three paradigms: Text-based (Fatemi et al., 2024; Luo et al., 2024; Chen et al., 2024; Guo et al., 2023; Wang et al., 2024) approaches leverage chain-of-thought prompting to decompose graph algorithms stepwise, but suffer error accumulation in multi-step reasoning. Tool-based (Zhang, 2023; Wang et al., 2025; OpenAI, 2023) methods offload computations to external solvers (e.g., NetworkX), yet rigid input requirements limit adaptability to diverse text representations. Agent-based (Li et al., 2024; Zhang et al., 2024) frameworks attempt task decomposition through multi-agent collaboration, but face scalability bottlenecks when handling large graphs due to code generation complexity and memory overload.
Working Memory Constraints in LLMs. Like humans facing working memory constraints in complex tasks (Jaeggi et al., 2008; O’Donohue et al., 2003), LLMs show similar capacity constraints. Cognitive studies show humans typically retain only few information units simultaneously, with performance decaying sharply in N-back tests (where N = 3) (Kirchner, 2006; Klatzky et al., 2008; Amon and Bertenthal, 2018; Jaeggi et al., 2010). LLMs mirror this pattern, exhibiting accuracy declines on 3-step dependencies (Baddeley and Hitch, 1974). This working memory overload manifests in graph reasoning as excessive graph scale and complex reasoning tasks, thus hindering effective processing of intricate graph structures and multi-step reasoning.
Graph Reasoning Benchmarks for LLMs. NLGraph (Wang et al., 2023), as the pioneering benchmark, demonstrates LLMs’ competence on small-scale graphs (less than 40 nodes) with basic reasoning tasks. GraphWiz (Chen et al., 2024) later expands to 100 nodes with more complex tasks. However, these and other benchmarks (Luo et al., 2024; Wang et al., 2024; Fatemi et al., 2024; Wang et al., 2025) remain constrained by fixed-size graphs within context windows and templated reasoning tasks, failing to match real-world reasoning demands where both scale and complexity exceed current test conditions.
3. Graph4real Construction
@@
Real-World Graph Scaling. To establish a comprehensive benchmark for evaluating LLMs on real-world graph reasoning tasks, we curate a diverse set of graphs from four domains: Web (Google Web Graph (McAuley and Leskovec, 2012)), Social (SNAP Social Circles (McAuley and Leskovec, 2012)), Transportation (PeMS (Shuman et al., 2013)), and Citation graphs (Cora (Craven et al., 1998)). Unlike prior benchmarks (Wang et al., 2023; Luo et al., 2024) that rely on randomly generated graphs, our dataset is constructed from real-world sources to ensure realistic topological properties and domain-specific characteristics. We employ a biased random walk sampling strategy to construct graphs at three scales (40, 100, and 1000 nodes), ensuring both effective evaluation of LLM-based methods through smaller-scale graphs and construct large-scale graphs to challenge existing methods.
@@
Figure 2. Graph N-back Query. Accuracy (right y-axis) measures memory retention across dialogues, computed as sum of True edges (correctly identified existing edges) and False edges (correctly rejected non-existent edges) (left y-axis).
@@
Figure 3. Graph Efficiency Coefficient (right y-axis) across sampling granularity. Edit shows graph Edit distance (left y-axis), and Cost (left y-axis, scaled ×100) counts total LLM interactions. Lower GEC indicates better performance.
Intent-Driven Task Design. A total number of 21 tasks are designed and categorized into three classes: (1) Graph Structural Querying Tasks: Focus on fundamental graph properties, such as edge existence or node count, testing basic graph structural comprehension. (2) Graph Algorithmic Reasoning Tasks: Require reasoning based on classical algorithms, such as shortest path and maximum flow, evaluating algorithmic proficiency. (3) Graph Predictive Modeling Tasks: Involve neural network-based predictions, such as node classification and link prediction, assessing predictive capabilities.
For task generation, we collect five real-world scenarios for each of the four domains (e.g., travel planning and logistics optimization for transportation) and craft 20 prompt templates to construct contextually grounded questions. Using DeepSeek-R1 (DeepSeek-AI, 2025), we generate graph reasoning tasks based on templates with specific scenarios, ensuring both diversity and relevance. To validate task quality, we employ a dual evaluation approach: DeepSeek-R1 for automated assessment and human for manual review. This process produces a dataset with 4200 questions distributed across the 21 tasks. The detailed task definitions, dataset statistics, and prompt templates are provided in our released anonymous github code.
4. GraphCogent
We propose GraphCogent, an agent-based framework inspired by the human working memory model to address the challenges of real-world graph reasoning tasks. Our framework consists of three specialized modules: Sensory Module (samples subgraphs and transform unstructured graph text representations into standardized adjacency list), Buffer Module (integrates graph data from Sensory Module and establishes data indices for diverse data formats), and Execution Module (combines tool calling and tool creation to address graph reasoning tasks). The overall pipeline of the GraphCogent is illustrated in Fig. 4.
4.1. Graph N-back Query
Real-world graphs present two fundamental characteristics: (1) Large Graph Scale, benchmarks like NLGraph (Wang et al., 2023) use small graphs (less than 40 nodes), while real-world graph (e.g., Cora (Craven et al., 1998)) contains thousands of nodes and edges; (2) Varied Text Representation, different domains use distinct descriptions ("connected" in transportation or "followed" in social), and same text representation may appear in varied formats (adjacency lists, symbolic notations, linguistic descriptions).
Figure 4. Overview of GraphCogent. Sensory Module (left) standardizes various graph text representations through subgraph sampling and conversion; Buffer Module (center) establishes cross-format data (e.g., NetworkX) integrating and indexing transformations; Execution Module (right) enables two reasoning modes: Execution Agent is employed for tool discrimination and implements tool calling for in-toolset tasks, Tool Creator handles out-toolset tasks based on tool creation.
@@ 4.1. Graph N-back Query
We propose Graph N-back Query, a graph memory evaluation for LLM inspired by the human working memory assessment through N-back tasks (Kirchner, 2006): Using a 100-node graph (sampled from PeMS datasets (Shuman et al., 2013)) partitioned into 50-edge subsets $E_{t}$, we test memory retention across three text representations by querying edge existence at turn $t+N$. Experiments with Llama3.1-8B in Fig. 3 reveal significant forgetting across dialogues, demonstrating that LLMs’ inherent memory mechanisms fail to reliably handle large-scale graph reasoning. This highlights that LLMs’ inherent context window limitations constrains their ability to maintain global graph information, rendering conventional step-by-step graph input approaches ineffective due to these working memory constraints.
4.2. Sensory Module